OpenStudy (anonymous):

The mean of 4 numbers is 18. What would the new mean be if the number 12 is added to the data set? A. 21 B. 30 C. 12.8 D. 16.8

OpenStudy (anonymous):

D

OpenStudy (anonymous):

sum X/ 4= 18 18 x 4 = X= 72 (72+ 12)/ 5= 16.8